Please help solving this rational equation:
2- (x-1)/(x+3)=(x-3)/(x-4)

Answer by algebrapro18(249)

2 - (x - 1)/(x + 3) = (x - 3)/(x - 4)
(2 - (x - 1)/(x + 3) = (x - 3)/(x - 4))(x + 3)(x - 4)
2(x - 3)(x - 4) - (x - 1)(x - 4) = (x - 3)(x + 3)
2(x^2 - 4x - 3x + 12) - (x^2 - 4x - x + 4) = x^2 - 9
2(x^2 - 7x + 12) - (x^2 - 5x + 4) = x^2 - 9
2x^2 - 14x + 24 - x^2 + 5x - 4 = x^2 - 9
x^2 - 9x + 20 = x^2 - 9
-9x + 29 = 0
-9x = -29
x = 29/9